OpenStudy (anonymous):

5 3/8 + 7 5/12 12 3/5 + 92 1/2 + 63 3/4

OpenStudy (lasttccasey):

\[5\frac{ 3 }{ 8 }+7\frac{ 5 }{ 12 }=12\frac{ 19 }{ 24 } or \frac{ 307 }{ 24 }\]

OpenStudy (lasttccasey):

\[12\frac{ 3 }{ 5 } + 92\frac{1}{2} + 63\frac{3}{4} = 168\frac{17}{20} or \frac{3377}{20}\]